OpenStudy (anonymous):

The of a fundraising event charges a $100 set-up fee and a $10 for each dinner.teh banquet hall for the event costs $1,500 to rent. If each person attending the event pays $50 for dinner, how many people must attend for the fundraising event to break-even?

OpenStudy (anonymous):

The Caterer*

OpenStudy (pfenn1):

To set up these equations you need to take stock of what information you are given, what the relationship is between the types of information and what are you trying to calculate. In this example, you want to find out how many people must attend to break even on the cost of the event. In other words, the number of people who attend, x, multiplied by the price they each pay to attend ($50 apiece) must equal the total cost that the event costs to run. The total cost will be the sum of the set-up fee ($100) plus the banquet hall rental fee ($1500) plus the cost of the dinner (number of dinners, x, times the cost per dinner, $10). So, can you write all this in the form of an equation?

OpenStudy (anonymous):

ahahahahaha

OpenStudy (anonymous):

no

OpenStudy (pfenn1):

The number of people who attend, x, multiplied by the price ($50 apiece) must equal the the sum of the set-up fee ($100) plus the banquet hall rental fee ($1500) plus the cost of the dinners (number of dinners, x, times the cost per dinner, $10). Try to translate these words into a mathematical equation. For instance, "the number of people, x, multiplied by the price ($50 apiece)" can be translated into x(50). How would you translate "the cost of the dinners (number of dinners, x, times the cost per dinner, $10)"?

OpenStudy (anonymous):

x=50+100+1500+10x

OpenStudy (anonymous):

rite?

OpenStudy (pfenn1):

In the following sentence, the resultant mathematical expression of each phrase I have put into the square brackets ( [ ] ). The number of people who attend, x, multiplied by the price ($50 apiece) [ 50x ] must equal [ = }the the sum of the set-up fee ($100) [ 100 ] plus [ + ] the banquet hall rental fee ($1500) [ 1500 ] plus [ + ] the cost of the dinners (number of dinners, x, times the cost per dinner, $10) [10x ]. The final equation is given by putting all the info in the square brackets together\[50x=100+1500+10x\] Do you see where I got each part of this equation?

OpenStudy (anonymous):

ya i was closle

OpenStudy (anonymous):

now what

OpenStudy (pfenn1):

Okay, now you need to rearrange the equation to solve for x. Can you do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

50x+1500=100+10x?

OpenStudy (pfenn1):

\[50x=1600+10x\]\[40x=1600\]Can you solve it from here?

OpenStudy (anonymous):

how is it 40 now?

OpenStudy (anonymous):

x= 40 but how did we get 40x from 50x

OpenStudy (pfenn1):

I started with 50x=1600+10x and subtracted 10x from both sides of the equation.\[50x-(10x)=1600+\cancel{10x}-\cancel{(10x)}\]Got it?

OpenStudy (anonymous):

yep x = 40

OpenStudy (pfenn1):

yep

OpenStudy (anonymous):

Lasto on..

OpenStudy (anonymous):

The force exerted by a 10-kg mass streches a spring 8 cm. how many centimeters(cm) will the spring strech if a 7-kg mas is used?

OpenStudy (pfenn1):

I don't remember my physics. What is the relationship between spring length and mass and force? You have to realize that I haven't studied physics in over 35 years.

OpenStudy (anonymous):

lol im only in chem

OpenStudy (anonymous):

F=ma F=mg F=kx mg=kx m/x=k/g m/x=m/x this is your ratio now go solve it

OpenStudy (anonymous):

tht dont rlly help meh but ok

OpenStudy (anonymous):

10/8=7/x solve x

OpenStudy (anonymous):

5.6 or \[X=\frac{ 28 }{ 5 }\]